# Dynamic Benchmark Construction for Evaluating Large Language Models on Real-World Codes

### v1

@@ H.2 Benchmark Task Examples
This appendix provides examples of representative benchmark instances from Code2Bench-2505. Each example showcases a complete task, including the instruction provided to the Large Language Model (LLM), the ground truth implementation from which the task was derived, the Property-Based Testing (PBT) script used for generating comprehensive test cases, and the test runner script for evaluating the LLM’s generated code.
H.2.1 SC Python Example: merge_json_recursive
@@
def merge_json_recursive(base, update):
"""Recursively merge two JSON-like objects.
[…]
- If both inputs are dictionaries, recursively merge them.
- If both inputs are lists, concatenate them.
- For all other cases, the update value overwrites the base value.
- The base object is left unmodified; a new merged object is returned.
[…]
base: Base JSON-like object (dictionary, list, or primitive value).
update: Update JSON-like object to merge into base.
[…]
A new JSON-like object containing merged content from base and update.
[…]
Input: base = {"a": 1}, update = {"a": 2}
Output: {"a": 2}
[…]
Output: [1, 2, 3, 4]
[…]
Output: {"a": {"b": 1, "c": 2}}
# TODO: Implement this function
pass
@@
from hypothesis import settings, given, Verbosity, example
from hypothesis import strategies as st
import json
import os
import atexit
import copy
[…]
TEST_CASE_DIR = os.path.abspath("test_cases")
os.makedirs(TEST_CASE_DIR, exist_ok=True)
TEST_CASE_FILE = os.path.join(TEST_CASE_DIR, "test_cases.json")
generated_cases = []
stop_collecting = False # Global flag to control case collection
[…]
def merge_json_recursive(base, update):
if not isinstance(base, dict) or not isinstance(update, dict):
if isinstance(base, list) and isinstance(update, list):
return base + update
return update
[…]
for key, value in update.items():
if key in merged:
merged[key] = merge_json_recursive(merged[key], value)
else:
merged[key] = value
[…]
json_strategy = st.recursive(
st.one_of([
st.integers(),
st.floats(allow_nan=False, allow_infinity=False),
st.text(st.characters(whitelist_categories=(’L’, N’, P’, S’, Z’))),
st.booleans()
lambda children: st.one_of(
st.lists(children, max_size=5),
st.dictionaries(st.text(st.characters(whitelist_categories=(’L’, N’)), max_size=5), children, max_size=5)
max_leaves=5
[…]
@settings(max_examples=10000, verbosity=Verbosity.verbose, print_blob=True)
@given(base=json_strategy, update=json_strategy)
def test_merge_json_recursive(base, update):
global stop_collecting
if stop_collecting:
return
[…]
update_copy = copy.deepcopy(update)
expected = merge_json_recursive(base_copy, update_copy)
[…]
generated_cases.append({
"Inputs": {"base": base, "update": update},
"Expected": expected
if len(generated_cases) >= 500:
stop_collecting = True
[…]
def save_test_cases():
with open(TEST_CASE_FILE, "w") as f:
json.dump(generated_cases, f, indent=2, ensure_ascii=False)
print(f"✅ Saved {len(generated_cases)} test cases to {TEST_CASE_FILE}")
[…]
@@
import json
import os
from tested import merge_json_recursive as func1
from helper import deep_compare
[…]
TEST_CASE_DIR = os.path.abspath("test_cases")
TEST_CASE_JSON_PATH = os.path.join(TEST_CASE_DIR, "test_cases.json")
[…]
if not os.path.exists(TEST_CASE_JSON_PATH):
print(f"JSON file not found: {TEST_CASE_JSON_PATH}")
return []
[…]
with open(TEST_CASE_JSON_PATH, "r") as f:
test_cases = json.load(f)
[…]
for i, case in enumerate(test_cases):
inputs = case["Inputs"]
expected_output = case["Expected"]
[…]
@@
actual_output = func1(**inputs) # Copy matrix to avoid in-place modification
[…]
if not deep_compare(actual_output, expected_output, tolerance=1e-6):
print(f"Test case {i + 1} failed:")
print(f" Inputs: {inputs}")
print(f" Expected: {expected_output}")
print(f" Actual: {actual_output}")
else:
print(f"Test case {i + 1} passed.")
[…]
test_cases = load_test_cases_from_json()
run_tests_with_loaded_cases(test_cases)
H.2.2 SC JavaScript Example: merge_json_recursive
This example demonstrates a Self-Contained (SC) task in JavaScript, requiring the recursive merging of JSON-like objects without external dependencies beyond standard library features.
Task Instruction
* Recursively merge two JSON-like objects.
* The function merges nested structures with the following rules:
* - If both inputs are dictionaries, recursively merge them.
* - If both inputs are lists, concatenate them.
* - For all other cases, the update value overwrites the base value.
* - The base object is left unmodified; a new merged object is returned.
* @param {any} base - Base JSON-like object (dictionary, list, or primitive value).
* @param {any} update - Update JSON-like object to merge into base.
* @returns {any} A new JSON-like object containing merged content from base and update.
* Examples:
* Input: base = {"a": 1}, update = {"a": 2}
* Output: {"a": 2}
* Input: base = [1, 2], update = [3, 4]
* Output: [1, 2, 3, 4]
* Input: base = {"a": {"b": 1}}, update = {"a": {"c": 2}}
* Output: {"a": {"b": 1, "c": 2}}
[…]
// TODO: Implement
Testcase Runner
import fs from fs’;
import path from path’;
import { merge_json_recursive as func } from ’./tested’;
[…]
const rawData = fs.readFileSync(filePath, utf-8’);
const testCases = JSON.parse(rawData);
return testCases;
[…]
const testCases = loadTestCases(’./3/test_cases/test_cases.json’);
[…]
test(‘Case {index}‘, () => {
let base_1 = tc.Inputs.base;
let update_1 = tc.Inputs.update;
[…]
return Math.abs(a - b) <= tolerance;
[…]
if (typeof a === "number" && typeof b === "number") {
return isClose(a, b, tolerance);
} else if (Array.isArray(a) && Array.isArray(b)) {
if (a.length !== b.length) return false;
return a.every((ai, index) => deepCompare(ai, b[index], tolerance));
} else if (
a !== null && typeof a === "object" &&
b !== null && typeof b === "object"
const aKeys = Object.keys(a);
const bKeys = Object.keys(b);
if (aKeys.length !== bKeys.length) return false;
return aKeys.every(k => k in b && deepCompare(a[k], b[k], tolerance));
} else {
return a === b;
H.2.3 SC TypeScript Example: merge_json_recursive
This example demonstrates a Self-Contained (SC) task in TypeScript, requiring the recursive merging of JSON-like objects without external dependencies beyond standard library features.
Task Instruction
* Recursively merge two JSON-like objects.
* The function merges nested structures with the following rules:
* - If both inputs are dictionaries, recursively merge them.
* - If both inputs are lists, concatenate them.
* - For all other cases, the update value overwrites the base value.
* - The base object is left unmodified; a new merged object is returned.
* @param base - Base JSON-like object (dictionary, list, or primitive value).
* @param update - Update JSON-like object to merge into base.
* @returns A new JSON-like object containing merged content from base and update.
* Examples:
* Input: base = {"a": 1}, update = {"a": 2}
* Output: {"a": 2}
* Input: base = [1, 2], update = [3, 4]
* Output: [1, 2, 3, 4]
* Input: base = {"a": {"b": 1}}, update = {"a": {"c": 2}}
* Output: {"a": {"b": 1, "c": 2}}
[…]
// TODO: Implement
Testcase Runner
import fs from fs’;
import path from path’;
import { merge_json_recursive as func } from ’./tested’;
[…]
const rawData = fs.readFileSync(filePath, utf-8’);
const testCases = JSON.parse(rawData);
return testCases;
[…]
const testCases = loadTestCases(’./3/test_cases/test_cases.json’);
[…]
test(‘Case {index}‘, () => {
let base_1: any = tc.Inputs.base;
let update_1: any = tc.Inputs.update;
[…]
return Math.abs(a - b) <= tolerance;
[…]
if (typeof a === "number" && typeof b === "number") {
return isClose(a, b, tolerance);
} else if (Array.isArray(a) && Array.isArray(b)) {
if (a.length !== b.length) return false;
return a.every((ai, index) => deepCompare(ai, b[index], tolerance));
} else if (
a !== null && typeof a === "object" &&
b !== null && typeof b === "object"
const aKeys = Object.keys(a);
const bKeys = Object.keys(b);
if (aKeys.length !== bKeys.length) return false;
return aKeys.every(k => k in b && deepCompare(a[k], b[k], tolerance));
} else {
return a === b;
H.2.4 SC Java Example: merge_json_recursive
This example demonstrates a Self-Contained (SC) task in Java, requiring the recursive merging of JSON-like objects without external dependencies beyond standard library features.
Task Instruction
package p3;
import java.util.Map;
import java.util.List;
[…]
* Recursively merge two JSON-like objects.
* The function merges nested structures with the following rules:
* - If both inputs are dictionaries, recursively merge them.
* - If both inputs are lists, concatenate them.
* - For all other cases, the update value overwrites the base value.
* - The base object is left unmodified; a new merged object is returned.
* @param base Base JSON-like object (dictionary, list, or primitive value).
* @param update Update JSON-like object to merge into base.
* @return A new JSON-like object containing merged content from base and update.
* Examples:
* Input: base = {"a": 1}, update = {"a": 2}
* Output: {"a": 2}
* Input: base = [1, 2], update = [3, 4]
* Output: [1, 2, 3, 4]
* Input: base = {"a": {"b": 1}}, update = {"a": {"c": 2}}
* Output: {"a": {"b": 1, "c": 2}}
public class Tested {
public static Object merge_json_recursive(Object base, Object update) {
// TODO: Implement this function
Testcase Runner
package p3;
[…]
import com.fasterxml.jackson.core.type.TypeReference;
import org.junit.jupiter.params.ParameterizedTest;
import org.junit.jupiter.params.provider.MethodSource;
import java.io.File;
import java.io.IOException;
import java.util.List;
import java.util.Map;
import java.util.stream.Stream;
import static org.junit.jupiter.api.Assertions.assertTrue;
[…]
// TestCase represents the structure of our test cases
static class TestCase {
public Inputs Inputs;
public Object Expected; // Use Object to support multiple types
[…]
public Object base; // Use Object to support Map or List
public Object update; // Use Object to support Map or List
[…]
private static List<TestCase> loadTestCases(String filePath) throws IOException {
ObjectMapper mapper = new ObjectMapper();
File file = new File(filePath);
try {
return mapper.readValue(file, new TypeReference<List<TestCase>>() {
} catch (IOException e) {
throw new IOException("Failed to load test cases from " + filePath + ": " + e.getMessage(), e);
[…]
private static Stream<TestCase> testCases() throws IOException {
return loadTestCases("src/test/java/test_cases/test_cases.json").stream();
[…]
@MethodSource("testCases")
void testMergeJsonRecursive(TestCase tc) {
try {
// Call the function under test
Object actual = Tested.merge_json_recursive(tc.Inputs.base, tc.Inputs.update);
[…]
assertTrue(
Helper.deepCompare(tc.Expected, actual, 0),
String.format(
"Test case failed:tc.Inputs.base, tc.Inputs.update, tc.Expected, actual));
} catch (Exception e) {
throw new RuntimeException("Unexpected exception during test execution", e);
H.2.5 SC Go Example: merge_json_recursive
This example demonstrates a Self-Contained (SC) task in Go, requiring the recursive merging of JSON-like objects without external dependencies beyond standard library features.
Task Instruction
package main
[…]
// The function merges nested structures with the following rules:
// - If both inputs are dictionaries, recursively merge them.
// - If both inputs are lists, concatenate them.
// - For all other cases, the update value overwrites the base value.
// - The base object is left unmodified; a new merged object is returned.
// Params:
// base: Base JSON-like object (dictionary, list, or primitive value).
// update: Update JSON-like object to merge into base.
// Returns:
// A new JSON-like object containing merged content from base and update.
// Examples:
// Input: base = {"a": 1}, update = {"a": 2}
// Output: {"a": 2}
// Input: base = [1, 2], update = [3, 4]
// Output: [1, 2, 3, 4]
// Input: base = {"a": {"b": 1}}, update = {"a": {"c": 2}}
// Output: {"a": {"b": 1, "c": 2}}
func merge_json_recursive(base, update interface{}) interface{} {
// TODO: Implement
Testcase Runner
package main
[…]
"encoding/json"
"fmt"
"os"
"testing"
[…]
type TestCase struct {
Inputs struct {
Base interface{} json:"base"‘
Update interface{} json:"update"‘
} json:"Inputs"‘
Expected interface{} json:"Expected"‘
[…]
func loadTestCases(filePath string) ([]TestCase, error) {
// Read the test cases file
file, err := os.ReadFile(filePath)
if err != nil {
return nil, fmt.Errorf("failed to read test cases: }
[…]
var testCases []TestCase
if err := json.Unmarshal(file, &testCases); err != nil {
return nil, fmt.Errorf("failed to parse test cases: }
[…]
func TestMergeJSONRecursive(t *testing.T) {
// Load test cases from the JSON file
testCases, err := loadTestCases("test_cases/test_cases.json")
if err != nil {
t.Fatalf("Failed to load test cases: }
[…]
for i, tc := range testCases {
t.Run(fmt.Sprintf("Case// Call the function under test
actual := merge_json_recursive(tc.Inputs.Base, tc.Inputs.Update)
[…]
if !DeepCompare(actual, tc.Expected, 0) {
t.Errorf(
Test case Input: base=Expected: Actual: i, tc.Inputs.Base, tc.Inputs.Update, tc.Expected, actual,
H.3 WSC Python Example: calculate_ngram_repetition
@@
from collections import Counter

### v2

@@ H.4 Benchmark Task Examples
This appendix provides examples of representative benchmark instances from Code2Bench-2509. Each example showcases a complete task, including the instruction provided to the Large Language Model (LLM), the ground truth implementation from which the task was derived, the Property-Based Testing (PBT) script used for generating comprehensive test cases, and the test runner script for evaluating the LLM’s generated code.
H.4.1 SC Python Example: merge_json_recursive
@@
1def merge_json_recursive(base, update):
2 """Recursively merge two JSON-like objects.
3
4 The function merges nested structures with the following rules:
5 - If both inputs are dictionaries, recursively merge them.
6 - If both inputs are lists, concatenate them.
7 - For all other cases, the update value overwrites the base value.
8 - The base object is left unmodified; a new merged object is returned.
9
10 Args:
11 base: Base JSON-like object (dictionary, list, or primitive value).
12 update: Update JSON-like object to merge into base.
13
14 Returns:
15 A new JSON-like object containing merged content from base and update.
16
17 Examples:
18 Input: base = {"a": 1}, update = {"a": 2}
19 Output: {"a": 2}
20
21 Input: base = [1, 2], update = [3, 4]
22 Output: [1, 2, 3, 4]
23
24 Input: base = {"a": {"b": 1}}, update = {"a": {"c": 2}}
25 Output: {"a": {"b": 1, "c": 2}}
26 """
27 # TODO: Implement this function
28 pass
@@ 25 Output: {"a": {"b": 1, "c": 2}}
1from hypothesis import settings, given, Verbosity, example
2from hypothesis import strategies as st
3import json
4import os
5import atexit
6import copy
7
8# Configuration
9TEST_CASE_DIR = os.path.abspath("test_cases")
10os.makedirs(TEST_CASE_DIR, exist_ok=True)
11TEST_CASE_FILE = os.path.join(TEST_CASE_DIR, "test_cases.json")
12generated_cases = []
13stop_collecting = False # Global flag to control case collection
14
15# Ground truth function
16def merge_json_recursive(base, update):
17 if not isinstance(base, dict) or not isinstance(update, dict):
18 if isinstance(base, list) and isinstance(update, list):
19 return base + update
20 return update
21
22 merged = base.copy()
23 for key, value in update.items():
24 if key in merged:
25 merged[key] = merge_json_recursive(merged[key], value)
26 else:
27 merged[key] = value
28
29 return merged
30
31# Strategy for JSON-like objects
32json_strategy = st.recursive(
33 st.one_of([
34 st.integers(),
35 st.floats(allow_nan=False, allow_infinity=False),
36 st.text(st.characters(whitelist_categories=('L', 'N', 'P', 'S', 'Z'))),
37 st.booleans()
38 ]),
39 lambda children: st.one_of(
40 st.lists(children, max_size=5),
41 st.dictionaries(st.text(st.characters(whitelist_categories=('L', 'N')), max_size=5), children, max_size=5)
42 ),
43 max_leaves=5
44)
45
46# Hypothesis test configuration
47@settings(max_examples=10000, verbosity=Verbosity.verbose, print_blob=True)
48@given(base=json_strategy, update=json_strategy)
49def test_merge_json_recursive(base, update):
50 global stop_collecting
51 if stop_collecting:
52 return
53
54 base_copy = copy.deepcopy(base)
55 update_copy = copy.deepcopy(update)
56 expected = merge_json_recursive(base_copy, update_copy)
57
58 if isinstance(base, (dict, list)) or isinstance(update, (dict, list)):
59 generated_cases.append({
60 "Inputs": {"base": base, "update": update},
61 "Expected": expected
62 })
63 if len(generated_cases) >= 500:
64 stop_collecting = True
65
66# Save test cases
67def save_test_cases():
68 with open(TEST_CASE_FILE, "w") as f:
69 json.dump(generated_cases, f, indent=2, ensure_ascii=False)
70 print(f" Saved {len(generated_cases)} test cases to {TEST_CASE_FILE}")
71
72atexit.register(save_test_cases)
@@
1import json
2import os
3from tested import merge_json_recursive as func1
4from helper import deep_compare
5
6# Configure save path
7TEST_CASE_DIR = os.path.abspath("test_cases")
8TEST_CASE_JSON_PATH = os.path.join(TEST_CASE_DIR, "test_cases.json")
9
10def load_test_cases_from_json():
11 if not os.path.exists(TEST_CASE_JSON_PATH):
12 print(f"JSON file not found: {TEST_CASE_JSON_PATH}")
13 return []
14
15 # Read JSON file
16 with open(TEST_CASE_JSON_PATH, "r") as f:
17 test_cases = json.load(f)
18
19 return test_cases
20
21def run_tests_with_loaded_cases(test_cases):
22 for i, case in enumerate(test_cases):
23 inputs = case["Inputs"]
24 expected_output = case["Expected"]
25
26 # Run function under test
@@
27 actual_output = func1(**inputs) # Copy matrix to avoid in-place modification
28
29 # Check if results match using deep_compare
30 if not deep_compare(actual_output, expected_output, tolerance=1e-6):
31 print(f"Test case {i + 1} failed:")
32 print(f" Inputs: {inputs}")
33 print(f" Expected: {expected_output}")
34 print(f" Actual: {actual_output}")
35 else:
36 print(f"Test case {i + 1} passed.")
37
38if __name__ == "__main__":
39 test_cases = load_test_cases_from_json()
40 run_tests_with_loaded_cases(test_cases)
H.5 WSC Python Example: calculate_ngram_repetition
@@
1from collections import Counter
2
3def calculate_ngram_repetition(text: str, n: int) -> float:
4 """
5 Calculates the proportion of repeated n-grams in a given text.
6
7 This function splits the input text into words and generates n-grams of the specified size `n`. It then computes the frequency of each n-gram and determines the proportion of n-grams that appear more than once. If there are no n-grams (e.g., when the text is empty or `n` is larger than the number of words in the text), the function returns 0.
8
9 Args:
10 text (str): The input text to analyze, consisting of words separated by spaces.
11 n (int): The size of the n-grams to generate (e.g., 2 for bigrams, 3 for trigrams).
12
13 Returns:
14 float: The proportion of n-grams that are repeated in the text. Returns 0 if no n-grams can be generated.
15
16 Raises:
17 ValueError: If `n` is less than or equal to 0.
18 """
19 # TODO: Implement this function
20 pass
Testcase Generator
1from hypothesis import settings, given, Verbosity, example
2from hypothesis import strategies as st
3import json
4import os
5import atexit
6import copy
7from collections import Counter
8
9# Configuration
10TEST_CASE_DIR = os.path.abspath("test_cases")
11os.makedirs(TEST_CASE_DIR, exist_ok=True)
12TEST_CASE_FILE = os.path.join(TEST_CASE_DIR, "test_cases.json")
13generated_cases = []
14stop_collecting = False # Global flag to control case collection
15
16# Ground truth function
17def calculate_ngram_repetition(text, n):
18 words = text.split()
19 ngrams = [tuple(words[i : i + n]) for i in range(len(words) - n + 1)]
20 ngram_counts = Counter(ngrams)
21 total_ngrams = len(ngrams)
22 repeated_ngrams = sum(1 for count in ngram_counts.values() if count > 1)
23 return repeated_ngrams / total_ngrams if total_ngrams > 0 else 0
24
25# Strategies for generating inputs
26def text_strategy():
27 return st.text(
28 alphabet=st.characters(whitelist_categories=('L', 'N', 'Z'), min_codepoint=32, max_codepoint=126),
29 min_size=0, max_size=100
30 )
31
32def n_strategy():
33 return st.integers(min_value=1, max_value=5)
34
35# Hypothesis test configuration
36@settings(max_examples=10000, verbosity=Verbosity.verbose, print_blob=True)
37@given(text=text_strategy(), n=n_strategy())
38@example(text="", n=1)
39@example(text="a", n=1)
40@example(text="a b c", n=2)
41@example(text="a a b b c c", n=2)
42@example(text="a b c d e f", n=3)
43@example(text="a a a a a a", n=3)
44def test_calculate_ngram_repetition(text, n):
45 global stop_collecting
46 if stop_collecting:
47 return
48
49 # Deep copy inputs to avoid modification
50 text_copy = copy.deepcopy(text)
51 n_copy = copy.deepcopy(n)
52
53 # Call func0 to verify input validity
54 try:
55 expected = calculate_ngram_repetition(text_copy, n_copy)
56 except Exception:
57 return # Skip inputs that cause exceptions
58
59 # Store inputs only
60 generated_cases.append({
61 "Inputs": {
62 "text": text_copy,
63 "n": n_copy
64 }
65 })
66
67 # Stop collecting after 500 cases
68 if len(generated_cases) >= 500:
69 stop_collecting = True
70
71# Save test cases
72def save_test_cases():
73 with open(TEST_CASE_FILE, "w") as f:
74 json.dump(generated_cases, f, indent=2, ensure_ascii=False)
75 print(f" Saved {len(generated_cases)} test cases to {TEST_CASE_FILE}")
76
77atexit.register(save_test_cases)
Testcase Runner
1import json
2import os
3import copy
4from collections import Counter
5from helper import deep_compare
6from tested import calculate_ngram_repetition as func1
7
8# Configure save path
9TEST_CASE_DIR = os.path.abspath("test_cases")
10TEST_CASE_JSON_PATH = os.path.join(TEST_CASE_DIR, "test_cases.json")
11
12# Ground truth function
13def calculate_ngram_repetition(text, n):
14 words = text.split()
15 ngrams = [tuple(words[i : i + n]) for i in range(len(words) - n + 1)]
16 ngram_counts = Counter(ngrams)
17 total_ngrams = len(ngrams)
18 repeated_ngrams = sum(1 for count in ngram_counts.values() if count > 1)
19 return repeated_ngrams / total_ngrams if total_ngrams > 0 else 0
20
21def load_test_cases_from_json():
22 if not os.path.exists(TEST_CASE_JSON_PATH):
23 print(f"JSON file not found: {TEST_CASE_JSON_PATH}")
24 return []
25 with open(TEST_CASE_JSON_PATH, "r") as f:
26 test_cases = json.load(f)
27 return test_cases
28
29def compare_outputs(expected, actual):
30 # Use deep_compare for basic types (int, float, str, etc.)
31 return deep_compare(expected, actual, tolerance=1e-6)
32
33def run_tests_with_loaded_cases(test_cases):
34 for i, case in enumerate(test_cases):
35 inputs = copy.deepcopy(case["Inputs"])
36 text = inputs["text"]
37 n = inputs["n"]
38
39 # Run ground truth and function under test
40 expected_output = calculate_ngram_repetition(text, n)
41 actual_output = func1(text, n)
42
43 # Compare outputs
44 if compare_outputs(expected_output, actual_output):
45 print(f"Test case {i + 1} passed.")
46 else:
47 print(f"Test case {i + 1} failed:")
48 print(f" Inputs: {inputs}")
49 print(f" Expected: {expected_output}")
50 print(f" Actual: {actual_output}")
51
52if __name__ == "__main__":
53 test_cases = load_test_cases_from_json()
54 run_tests_with_loaded_cases(test_cases)
Appendix I Detailed Evaluation
I.1 Compute Resources
we detail the compute resources utilized for evaluating the Large Language Models on the Code2Bench-2509 benchmark. The evaluation was conducted on an infrastructure consisting of server-grade machines. Open-source models with fewer than 32B parameters (as listed in Table 2) were served using vLLM on a cluster equipped with NVIDIA GPUs. Specifically, these models were evaluated on machines featuring NVIDIA A100 80GB GPUs. The evaluation environment for these models was containerized to maintain isolation and consistency. For larger open-source models (>= 32B parameters) and all closed-source models, evaluation was performed by accessing their respective official APIs. The compute resources for these API-based evaluations are managed by the model providers and are not under our direct control or knowledge. Therefore, we cannot provide specific details on the underlying hardware, memory, or parallelization used by these providers. The Testcase Runner execution for each task (which involves loading test cases, running the generated code and ground truth, and performing differential testing) was primarily CPU-bound and ran on standard server CPUs, such as Intel(R) Xeon(R) Gold 5218 CPU @ 2.30GHz, featuring 64 logical cores. These machines were equipped with 125 GiB of RAM and SSD storage for the benchmark data and test cases. The total compute time required for the comprehensive evaluation of all 16 models across the 1163 tasks in Code2Bench-2509 was substantial. While precise timing varies per model and task, we estimate the total GPU-hours consumed for the open-source model inferences to be approximately 200 GPU-hours. The total CPU-hours consumed for Testcase Runner execution across all models (including running the generated code and ground truth against approximately 500 tests per task per model) is estimated to be approximately 200 CPU-hours.
Benchmark Construction Cost. The construction of the Code2Bench-2509 suite is a computationally intensive but one-time investment. Generating a rigorous Property-Based Testing (PBT) suite with a guaranteed 100% branch coverage takes approximately 5 CPU-minutes per task. This process includes the iterative generation of PBT driver code by the LLM, execution of test cases, and coverage verification. Constructing the entire suite required approximately 83 CPU-hours. Crucially, this pipeline is embarrassingly parallelizable, allowing the entire benchmark to be regenerated in under one hour on a standard high-performance computing cluster, making continuous dynamic updates highly practical.